Ethene is converted to ethane by the reaction

C2H4 flows into a catalytic reactor at 28.3 atm and 250.°C with a flow rate of 1150. L/min. Hydrogen at 28.3 atm and 250.°C flows into the reactor at a flow rate of 1400. L/min. If 14.0 kg C2H6 is collected per minute, what is the percent yield of the reaction?
1
Expert's answer
2017-11-27T11:37:06-0500
The reaction equation: C2H4 + H2 = C2H6
The amount of ethene used per minute can be find from the ideal gas law:
PV = nRT → n = PV / RT
Converting data to standard units (SI):
P = 28.3 atm = 28.3 ∙ 101325 Pa = 2867500 Pa
V = 1150 L = 1.15 m3
T = 250°C = 250 + 273 = 523 K
R = 8.314 J∙mol-1∙K-1 (universal gas constant)
So, the amount of C2H4 per minute:
n = (2867500 ∙ 1.15) / (8.314 ∙ 523) = 758.4 mol
The amount of C2H6 per minute:
n’ = 14000 / 30 = 466.7 mol
Therefore, the yield:
(n’ / n) ∙ 100% = 0.615 ∙ 100% = 61.5%
Answer:
61.5%